OpenStudy (anonymous):

Find the first six terms of the sequence. a1 = -8, an = 5 • an-1

OpenStudy (anonymous):

@ghazi

OpenStudy (ghazi):

\[A _{1}=-8, A _{n}=5A_{n-1}\] is that your question?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

\[a_2=5\times a_1=5\times (-8)=-40\] for the first one

OpenStudy (anonymous):

then \[a_3=5\times a_2=5\times (-40)=200\] for the next one and so on